OpenStudy (anonymous):

I understand everything but putting the equation into two binomials that when multiplied together equal that equation. I know how to do that with normal numbers, but not decimals. It got me really confused A juggler is performing an act by juggling several balls. The juggler throws the balls up at an initial height of 3.9 feet with a speed of 14.8 feet per second. If the juggler did not catch a ball, about how long will it take the ball to hit the floor? using this equation?H(t)=-16t^2+vt+s

OpenStudy (shane_b):

The kinematic equation is: \[h(t)=-16t^2+vt+s\] Given the values in the problem it becomes:\[h(t)=-16t^2+(14.8ft/s)t+3.9ft=0\]All you have to do is solve for t. You will get two values...one will obviously be 0 and the other will be the total flight time.

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20110819123824AAJlbIG h(t) = 0 = 3.9 + 14.8t - 16.1t² 16.1t² - 14.8t - 3.9 = 0 t = [14.8 ± √(14.8² + 4*16.1*3.9)]/32.2 ≈ 1.14 seconds (the t < 0 root can be discarded as outside the domain)

OpenStudy (shane_b):

Well, I guess one answer will actually be -.0214s...but as @Yahoo! said, you can disregard it since you can't have a negative time.